### v2

@@ 4.2 EVA results
Phikon-V2 307M 0.727 0.939 0.775 0.893 0.808 0.635 0.630 0.639 0.760
@@ 4.3 Robustness evaluation
Figure 1: PLISM results - Median Top-5 Accuracy vs. Median Cosine Similarity by extractor. For both axes, higher values indicate more robust models.
Figure 1 presents the robustness metrics evaluated on the PLISM dataset. H0-mini achieves notably high scores, outperforming state-of-the-art models. We also note that multi-modal extractors, such as CONCH, demonstrate high robustness. We note that while there is a correlation between the cosine similarity metric and the retrieval metric, cosine similarity alone is not enough to assess the robustness of a foundation model. Models, such as PLIP and Phikon, present high values of cosine similarity, while failing at the retrieval task. Besides, robustness of a model should be evaluated alongside downstream performance, to have a comprehensive understanding of a model’s behavior. It should also be noted that robustness to staining is harder to achieve compared to scanning as staining variations may have an impact on biological morphology (different stainings correspond to different consecutive cuts) while scanning variations mostly impact color tints. Figure 4 in Appendix 7.2 shows for all models their respective robustness with respect to scanning or staining variations based on Top-1 Accuracy. H0-Mini presents the highest robustness to scanning variations and ranks 3rd on staining robustness behind H-Optimus-0 and Virchow2.
@@ 5 Discussion
While the results presented in this work are very promising, several limitations and perspectives can be noted. First, we note that the final projection head of the teacher model is required to perform the distillation. Whether the distillation can be performed without the projection head (for instance, by learning them during the pre-training as in [23]) with equivalent downstream performance and robustness should be explored.
Second, even though distilled models will be much cheaper to use for downstream tasks (both in time, cost or energy consumption), it remains computationally demanding to obtain them compared to training a similar model from scratch (1.7x longer for a ViT-B architecture) as it requires the teacher model. However, the resulting performance and robustness is greatly improved.
Finally, we mention some open questions. The successful distillation of large models into smaller ones raises a question on the intrinsic dimension of foundation models in digital pathology, that should be further investigated. Besides, we did not perform an extensive hyperparameter optimization. Whether some could be tailored for distillation remains to be explored (e.g., adding more global crops to the iBOT objective, reducing the student patch size, etc.). The choice of dataset on which to perform distillation may also have an impact on the downstream properties of the models.
@@ 6 Conclusion
This study shows that distilling a large foundation model results in a competitive smaller model. We show that this model can efficiently be leveraged for downstream applications closing the performance gap with larger foundation models. We hope our work will democratize the use of pathology foundation models for researchers with limited computational resources.
@@
This work was granted access to the High-Performance Computing (HPC) resources of IDRIS under the allocations 2023-A0141012519 and 2024-A0161012519 made by GENCI.
@@
The results presented here are in part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga.

### v3

@@ 3.1 Performance evaluation
PhikonV2 307M 0.727 0.939 0.775 0.893 0.808 0.635 0.630 0.639 0.760
In Table 2, we report the results on the HEST benchmark. H0-mini also performs well on this benchmark highlighting the efficiency of knowledge distillation. The model significantly outperforms Phikon, its ViT-B equivalent trained from scratch with iBOT on the same pre-training data. Moreover, it surpasses on average much bigger foundation models such as Virchow2 or Gigapath.
Table 2: HEST results (Pearson correlations). For all models, [CLS] tokens are concatenated to the mean over all patch tokens to form the input embeddings.
Model Idc Prad Paad Skcm Coad Read Ccrcc Luad L. Idc Mean
UNI2-h 0.6054 0.3753 0.5231 0.6829 0.3319 0.2265 0.2662 0.5743 0.2743 0.4292
H0 0.6106 0.3621 0.5106 0.6614 0.3089 0.2401 0.2669 0.5754 0.2664 0.4224
H0-mini 0.5909 0.3633 0.5068 0.6125 0.2700 0.2047 0.2643 0.5633 0.2640 0.4044
Virchow2 0.5971 0.3528 0.4778 0.6404 0.2580 0.2073 0.2604 0.5685 0.2568 0.4019
Hibou L16 0.5945 0.3231 0.4758 0.6059 0.3128 0.1823 0.2777 0.5720 0.2490 0.3992
Gigapath 0.5707 0.3841 0.4920 0.5823 0.3076 0.186 0.2277 0.5579 0.2499 0.3952
GPFM 0.5796 0.3733 0.4686 0.5839 0.2801 0.1769 0.2510 0.5522 0.2391 0.3940
Kaiko B8 0.5710 0.3827 0.4727 0.5904 0.3105 0.1726 0.2664 0.5883 0.2362 0.3912
UNI 0.5851 0.3274 0.4882 0.6235 0.2583 0.1757 0.2463 0.5558 0.2576 0.3907
PhikonV2 0.5677 0.3793 0.4771 0.5845 0.2561 0.1865 0.2607 0.5502 0.2476 0.3897
Phikon 0.5481 0.3452 0.4639 0.5555 0.2668 0.1667 0.2496 0.5679 0.2387 0.3780
Overall ranking and comparison between models is performed through a one-sided Wilcoxon sign rank test with Holm correction (at $\alpha=0.05$) on HEST and EVA combined (17 tasks). This places H0-mini at 4th position, outperformed by UNI2-h ($p<10^{-4}$), H-Optimus-0 ($p=0.039$) and Virchow2 (n.s., $p=0.29$). H0-mini statistically outperforms UNI ($p=0.044$) and the Phikon models ($p<10^{-3}$).
@@ 3.2 Robustness evaluation
Table 3 presents the robustness metrics evaluated on the PLISM dataset. H0-mini achieves high scores, outperforming all other state-of-the-art models. We also note that multi-modal extractors, such as CONCH, demonstrate high robustness. While there is a correlation between the cosine similarity metric and the retrieval metric, cosine similarity alone is not enough to assess the robustness of a foundation model. Models such as PLIP [11] and Phikon present high values of cosine similarity, while failing at the retrieval task. It should also be noted that robustness to staining is harder to achieve compared to scanner as staining variations may have an impact on biological morphology while scanning variations mostly impact color tints. Looking at top-10 accuracy, H0-Mini presents the highest robustness to scanning variations, as well as joint scanning and staining variations.
Table 3: PLISM results (cosine similarity and top-10 accuracy). For each metric, we report median and inter-quartile range over the corresponding slide pairs.
Fixed-staining, cross-scanner Fixed-scanner, cross-staining Cross-staining, cross-scanner
Model Cosine top-10 Cosine top-10 Cosine top-10 $\downarrow$
H0-mini 0.92 (0.06) 0.86 (0.28) 0.83 (0.09) 0.32 (0.33) 0.79 (0.11) 0.18 (0.25)
H0 0.85 (0.08) 0.74 (0.32) 0.73 (0.11) 0.33 (0.29) 0.67 (0.12) 0.17 (0.20)
Virchow2 0.88 (0.06) 0.61 (0.31) 0.81 (0.09) 0.31 (0.25) 0.77 (0.10) 0.16 (0.18)
CONCH 0.93 (0.04) 0.75 (0.29) 0.86 (0.05) 0.24 (0.21) 0.84 (0.06) 0.16 (0.16)
GigaPath 0.79 (0.11) 0.59 (0.40) 0.61 (0.13) 0.12 (0.18) 0.56 (0.14) 0.05 (0.08)
UNI 0.76 (0.09) 0.53 (0.30) 0.62 (0.13) 0.17 (0.22) 0.53 (0.13) 0.05 (0.08)
UNI2-h 0.76 (0.17) 0.50 (0.56) 0.68 (0.13) 0.19 (0.25) 0.56 (0.17) 0.05 (0.10)
Kaiko B8 0.87 (0.07) 0.35 (0.46) 0.81 (0.09) 0.15 (0.20) 0.75 (0.11) 0.05 (0.09)
GPFM 0.80 (0.08) 0.36 (0.34) 0.69 (0.17) 0.09 (0.18) 0.58 (0.14) 0.02 (0.04)
Hibou L16 0.65 (0.10) 0.06 (0.09) 0.61 (0.16) 0.03 (0.06) 0.47 (0.14) 0.008 (0.01)
PLIP 0.92 (0.04) 0.05 (0.20) 0.91 (0.05) 0.04 (0.10) 0.87 (0.06) 0.004 (0.01)
Phikon 0.81 (0.08) 0.13 (0.24) 0.69 (0.19) 0.02 (0.07) 0.60 (0.18) 0.004 (0.01)
PhikonV2 0.72 (0.09) 0.06 (0.10) 0.66 (0.18) 0.03 (0.07) 0.54 (0.14) 0.003 (0.008)
Finally, table 4 presents the downstream results on the BreastBm dataset. H0-mini consistently and significantly outperforms H-Optimus-0 on each of the performance and robustness metrics ($p<10^{-4}$) as assessed by multiple boot-strap-based comparison tests combined with harmonic mean p-value across subcohorts (resp. subcohort pairs) for AUC (resp. CCC). Considering that the features aggregation module is a simple average, BreastBm results suggest that robustness in the feature space translates to downstream robustness.
Table 4: Performance (measured by AUC) and robustness (measured by CCC) results on the BreastBM dataset.
Performance metric: AUC Robustness metric: CCC
Task Model Clb Gr Curimeta Clb Gr Curimeta
gBrca H0-mini 0.69 (0.02) - 0.69 (0.01) 0.92 (0.03) - 0.88 (0.84-0.91)
H0 0.66 (0.02) - 0.64 (0.05) 0.42 (0.25) - 0.71 (0.65-0.78)
ER H0-mini 0.95 (0.00) 0.76 (0.02) 0.88 (0.00) 0.97 (0.02) 0.82 (0.08) 0.98 (0.97-0.98)
H0 0.95 (0.01) 0.75 (0.03) 0.88 (0.00) 0.97 (0.01) 0.72 (0.12) 0.90 (0.87-0.95)
PR H0-mini 0.89 (0.01) 0.87 (0.00) 0.69 (0.01) 0.96 (0.03) 0.78 (0.10) 0.98 (0.97-0.98)
H0 0.88 (0.01) 0.88 (0.02) 0.69 (0.01) 0.92 (0.03) 0.63 (0.18) 0.94 (0.92-0.96)
HER2 H0-mini - 0.66 (0.02) - - 0.74 (0.14) -
H0 - 0.65 (0.02) - - 0.64 (0.18) -
@@ 4 Discussion
While the results presented in this work are promising, several limitations and perspectives can be noted. First, we note that the final projection head of the teacher model is required to perform the distillation. Whether the distillation can be performed without the projection head (for instance, by learning them during the pre-training as in [17]) with equivalent downstream performance and robustness should be explored. Second, even though distilled models will be much cheaper to use for downstream tasks (both in time, cost or carbon footprint), they remain computationally demanding to train compared to training a similar model from scratch (1.7x longer for a ViT-Base) as it requires the teacher model. However, the resulting performance and robustness is greatly improved.
Finally, we mention some open questions. The successful distillation of large models into smaller ones raises a question on the intrinsic dimension of foundation models in digital pathology. Besides, we did not perform an extensive hyperparameter optimization. Whether some could be tailored for distillation remains to be explored (e.g., adding more global crops, reducing the student patch size, etc.). Additionally, we proposed several metrics to assess the robustness of foundation models, yet future work should be carried out to have a more comprehensive understanding of a model’s behavior [13]. Note that this may require high-quality multi-scanned, multi-stained datasets. We hope that releasing our processed version of PLISM will encourage further research in this direction.
@@ 5 Conclusion
This study shows that distilling a large foundation model results in a competitive smaller model. We show that this model can efficiently be leveraged for downstream applications closing the performance gap with larger foundation models. By releasing the model publicly, we hope our work will democratize the use of pathology foundation models for researchers with limited computational resources.
@@ 5.0.1 Acknowledgements
This work was granted access to the High-Performance Computing (HPC) resources of IDRIS under the allocations 2023-A0141012519, 2024-A0161012519 and 2024-GC011015442 made by GENCI.
@@
The results presented here are in part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga. Results on the gBRCA dataset are based on data generated by Gustave Roussy (GR) and Centre Léon Bérard (CLB) through PortrAIt (Owkin, Tribun Health, Cypath, GR, CLB, Unicancer), a French consortium advancing precision medicine. PortrAIt was funded by the French government as part of the France 2030 program, and the European Union – NextGenerationEU through the France Relance program An additional external data set from the United States was accessed through CuriMeta.
Appendix 0.A Supplemental
0.A.1 Distillation hyperparameters
We provide in Table 5 the main hyperparameters used in the distillation.
Table 5: Main hyperparameters used in the distillation.
Optimization Warmup epochs 16
Parameters Teacher temperature warmup epochs 30
Weight decay end value 0.4
Total batch size 2,048
Number of iterations 105,000
Model Patch size 14
Parameters Register tokens 4
Embedding dimension 768
Layers 12
Heads 12
MLP ratio 4
MLP activation SwiGLU
Projection Heads prototypes 131,072
Heads DINO head bottleneck dim 384
iBOT head bottleneck dim 256
Hardware GPUs 128 V100 32 Go
0.A.2 Mathematical details of robustness evaluation metrics
0.A.2.1 Cosine Similarity between Tiles
The cosine similarity between two tiles is a measure of the similarity between their feature vectors. Let $t_{1}$ and $t_{2}$ represent the feature vectors of two tiles. The cosine similarity between $t_{1}$ and $t_{2}$ is defined as:
Cosine $\text{Cosine Similarity}(t_{1},t_{2})=\frac{\langle t_{1},t_{2}\rangle}{\|t_{1}\|_{2}\|t_{2}\|_{2}},$
where
  • $\bullet$
[…]
  • $\bullet$
    and $\|\cdot\|_{2}$ denotes the corresponding euclidean norm.
0.A.2.2 Cosine Similarity for a Slide Pair
For each slide pair $(S_{1},S_{2})$, the Mean Cosine Similarity is computed by averaging the cosine similarities between corresponding tile pairs. Each tile pair $(t_{i,1},t_{i,2})$ corresponds to a matched location, where $t_{i,1}$ and $t_{i,2}$ are the feature vectors extracted from slides $S_{1}$ and $S_{2}$, respectively. Let $N$ denote the total number of tile pairs. The Mean Cosine Similarity between $S_{1}$ and $S_{2}$ is then:
$\text{Mean Cosine Similarity}_{S_{1},S_{2}}=\frac{1}{N}\sum_{i=1}^{N}\text{Cosine Similarity}(t_{i,1},t_{i,2}).$
0.A.2.3 Top-k Accuracy for a Slide Pair
For each tile $t_{i,1}$ from slide $S_{1}$, the cosine similarities between $t_{i,1}$ and all other tiles from both $S_{1}$ and $S_{2}$ are computed. These tiles are then ranked based on their similarity. The Top-k Accuracy is defined as the fraction of tiles from $S_{1}$ whose corresponding tile from $S_{2}$ ranks among the top $k$ closest tiles. Formally, the Top-k Accuracy from $S_{1}$ to $S_{2}$ is:
[…]
where the rank of $t_{i,2}$ in $S_{1}\cup S_{2}$ is defined as:
$\text{rank}_{S_{1}\cup S_{2}}(t_{i,2})=\left|\{t\in(S_{1}\cup S_{2}),t\neq t_{i,1}:\text{Cosine Similarity}(t_{i,1},t)\geq\text{Cosine Similarity}(t_{i,1},t_{i,2})\}\right|.$
The final Top-k Accuracy for the slide pair is obtained by averaging the Top-k Accuracy in both directions:
[…]